OpenStudy (cicilybailey14):

Write the equation in standard form using integers: y = 7/3x + 25/3 -7x + 3y = 25 -3x + y =3 7x + y = -25 -7x + y = 25

OpenStudy (anonymous):

? ope = 4.667/2.000 = 2.333 ?smex -intercept = 25/-7 = -3.57143 ?hi -intercept = 25/3 = 8.33333 they wont let meh give u direct answers sooo XD

OpenStudy (cicilybailey14):

I got it thanks :*)

OpenStudy (anonymous):

HA! np bruh

OpenStudy (anonymous):

get rid of the fractions, multiply the entire equation by 3 because the denominator is 3 3y = 7x + 25 move the x & y to one side and you get -7x + 3y = 25

OpenStudy (cicilybailey14):

Thats what I got too @kevinngo :D

OpenStudy (anonymous):

glad you got the same one